OpenStudy (anonymous):

Establish identity: what is sec(pheta)sin(pheta)/tan(pheta)-1?

OpenStudy (anonymous):

sec=1/cos tan=sin/cos

Directrix (directrix):

Is this the expression you posted? See attachment. @jkfussball

OpenStudy (anonymous):

the answer is zero b/c secsin becomes tan. so tan-tan=0.

Directrix (directrix):

@jkfussball What about the 1 in the expression?

OpenStudy (anonymous):

that becomes tan/tan

OpenStudy (anonymous):

do you know what cos5pie/12 is using the difference identity?

OpenStudy (bradely):

Tan (theta) = sin(theta)/cos(theta) sec(theta)sin(theta)/tan(theta)-1 sec(theta)sin(theta)/[sin(theta)/cos(theta)]-1 cancel common terms sec(theta)/[1/cos(theta)]-1 1/cos(theta)=sec(theta) sec(theta)/sec(theta)-1 1-1 =0 Source: http://www.mathskey.com/question2answer/